OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

Ridiculous.

OpenStudy (anonymous):

We will say you are given the expression x^2+15x+56 and you are asked to find the other factor when the first factor you received is x+8 . long division x+8 into x^2+15x+56 always use the leading coefficient of what you want to go into something such as x+8 into X^2+15x+56 how many times does x go into x^2 - x^2+8x -------------------- since you are dividing you would subtract to get your answer x^2-x^2=0 15x-(+8x) 7x + 56 is equal to 7x how many times does x go into 7x x+7 -7x+56= -7x-56 7x+56 -7x-56 ----------- 0 or you could set it up like a fraction which is really what division really is x^2+15x+56 if the polynomial factors here you can use this method --------------------- x+8 x^2+15x+56 (x+8) (x+7) what multiplied comes out to 56 and added will equal 15 8,7 (x+8)(x+7) ---------------- (x+8) the 2 x+8 's cancel and you have x+7 as your answer

OpenStudy (anonymous):

(3x^3 - 6x^2 + x) / 3x^2 3x^3/3x^2 - 6x^2 / 3x^2 + x/ 3x^2 x^1 - 2x^0 + 1/x^1 x - 2 +1/3x You divide the first term 3x^3 by 3x^2 and 3 divided by 3 gives you 1 and dividing exponents you subtract so 3 - 2 gives you the x^1 which is actually just x The second term is -6x^2 divided by 3x^2 and -6 divided by 3 is -2 and x^2 - x^2 is 0 and x^0 will give you one so -2 times 1 will be -2 The third term is x divided by 3x^2 so 1 - 2 is -1 so the x is on the bottom (denominator) with the 3 and you need a 1 on the top (numerator) 1/3x so if I take x-2 + 1/3x and I multiply by 3x^2 I get 3x^3 - 2x^2 + 3x^2/ 3x which simplifies to x which is what we started out with.

OpenStudy (ash2326):

Let's our polynomial be \[x^2+3x+3\] For simplicity I've taken a quadratic:)

OpenStudy (anonymous):

Okay.

OpenStudy (ash2326):

Sorry it's \[x^2+3x+2\]

OpenStudy (anonymous):

Alright.

OpenStudy (ash2326):

Let the monomial be \[x+1\]

OpenStudy (anonymous):

Long....

OpenStudy (ash2326):

We have to divide the two \[\frac{x^2+3x+2}{x+1}\] Let's first simplify this by long division \(x\) _________________________________ \(x+1\) | \(x^2+3x+2\) -(\(x^2+x\)) \(x+2\) Next step \(x+2\) _________________________________ \(x+1\) | \(x^2+3x+2\) -(\(x^2+x\)) \(2x+2\) -( \(2x+2\) ) ___________________________ 0 So we get \[\large \frac{x^2+3x+2}{x+1}=x+2\]

OpenStudy (ash2326):

@careless850 did you understand till here?

OpenStudy (anonymous):

It looks all deformed at the top..

OpenStudy (anonymous):

Confusing.

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

I can barely read it. O.O

OpenStudy (ash2326):

Sorry, do you understand this